# Advanced computer vision for extracting georeferenced vehicle trajectories from drone imagery11footnotemark: 1

### v2

@@ Abstract
Drones , Traffic monitoring , Machine vision , Vehicle tracking , Video image processing
@@ 1 Introduction
Traditional traffic monitoring methods, such as loop detectors and manual counting, are becoming obsolete in the era of smart cities due to their inflexibility and limited scope [1]. Stationary camera surveillance, although effective, presents economic challenges due to high installation costs and limited field of view (FoV), leading to blind spots. Data from global navigation satellite system (GNSS) and connected vehicles also lack the versatility needed for comprehensive traffic monitoring. Several vehicle detection and tracking technologies have emerged, including camera-based [2, 3, 4], LiDAR-based [5, 6], magnetics-based [7, 8], and radar-based [9, 10] systems. While each offers certain advantages, they still face limitations in adaptability and coverage. Accurate estimation of traffic states allows for adjustments in traffic control strategies [11], enhancing network utilization [12, 13], improving road user experiences, and yielding economic and ecological benefits.
@@ 2.5 Availability of datasets
Vehicle annotation datasets, such as CARPK [110], CyCAR [111], HARPY [112], RAI4VD [113], UAVDT [114], UIT-ADrone [115], and VisDrone [116], serve as valuable training resources for object detection models. Nonetheless, they frequently utilize basic annotation methods, cover a restricted or uncoherent range of vehicle classes, and exhibit varying degrees of annotation accuracy, thereby potentially affecting the reliability and consistency of DL models trained on these datasets. To mitigate these limitations, we incorporated these datasets in conjunction with our custom Songdo Vision dataset [29] and employed a multi-phase object detector training pipeline, which is described in Section 3.3, where Table 1 presents the relevant metadata for these datasets.
@@ 3.3 Object detection
The datasets used for training, comprising images and annotations from eight public datasets and our Songdo Vision dataset (detailed in Section 3.2), are summarized in Table 1, which also specifies their usage across the BASE and FINE training stages. The public datasets include a variety of aerial images, characterized by varying altitudes, camera angles, perspectives, resolutions, illumination conditions, image quality, and train/test split ratios. For the public datasets, we adhered to the original train/test splits whenever available. In cases where no such split was provided, we applied an 80/20 division to maintain uniformity. Furthermore, datasets containing a high proportion of repetitive images were proportionally downsampled to maintain a balanced representation across sources. This ensured that no single dataset dominated the training process, enhancing the model’s ability to generalize across diverse traffic conditions.
Table 1: Datasets employed in multi-stage training.
@@
To ensure consistency with our 140m–150m BEV scenes in Songdo, we performed extensive preprocessing and filtering of these datasets. Although a comprehensive description of these steps is omitted here for brevity, we focused on consolidating the diverse class labels found in the public datasets to streamline our trajectory extraction process, which utilizes four primary vehicle categories and two auxiliary categories: $\mathcal{C}={0,,1,,2,,3,,4,,5}$, where 0 corresponds to cars (including vans and light-duty vehicles), 1 to buses, 2 to trucks, 3 to motorcycles, 4 to pedestrians, and 5 to bicycles. While only the four primary vehicle categories were used during inference, the auxiliary categories (pedestrians and bicycles) were retained during training to improve overall classification performance.
To ensure that our FINE dataset meets stringent training specifications, we excluded certain datasets, such as VisDrone and UAVDT. For example, VisDrone’s annotations primarily originate from non-BEV environments at lower altitudes, resulting in inconsistent object size representations and varied camera settings that complicated integration. Additionally, this dataset exhibited mediocre annotation accuracy, further justifying its exclusion from our FINE dataset.
@@
Table 2: Annotation counts and class distribution across training and test sets for the considered datasets.
@@
Songdo Vision, comprising 5,419 annotated images, was split into training and test sets, with 20% (1,084 images) designated as the test set, allocated for both model validation during training and final testing of our top-performing object detector. Table 2 shows the class distribution across the sets. Our top-performing model was selected based on its highest mean Average Precision (mAP) at an IoU of 0.5 (mAP@50) across all classes. As shown in Table 3, this model demonstrated remarkable performance across all vehicle types, achieving an overall mAP@50 of 0.951 and an exceptional mAP@50 of 0.888 for motorcycles, despite their smaller pixel size. Beyond conventional detection metrics, we also assessed the positional accuracy of BB centers, revealing a mean Euclidean error of approximately 2.21 pixels (12.2 cm), further reinforcing the reliability of our detection step for trajectory extraction. These results highlight the model’s robustness in detecting small objects in BEV drone videos. To briefly define the key metrics: Precision indicates the proportion of correctly identified vehicles out of the total detected, while Recall reflects the ratio of correct detections to the actual number of vehicles in a class. The mAP@50 measures detection accuracy using an IoU threshold of 0.5, and mAP@50-95 evaluates performance over a range of IoU thresholds (0.5 to 0.95), with a step size of 0.05, capturing variability in BB overlaps.
@@
The confusion matrix in Fig. 4 (left) illustrates the model’s classification accuracy and misclassification patterns across vehicle classes. The precision-recall curve in Fig. 4 (right) further demonstrates strong detection performance, with low false positive and false negative rates. While smaller objects like pedestrians and bicycles are occasionally misclassified as motorcycles, particularly near pedestrian crossings or bike lanes, these instances were not filtered due to the inherent challenges in accurate differentiation. Despite this, the model’s robust performance on motorcycles, though slightly lower than for larger vehicles, highlights the challenges of detecting smaller objects in drone footage.
@@ 3.5.2 Method and parameter search campaign
We implemented several prominent local feature detectors including SIFT, RSIFT, BRISK, ORB, KAZE, and A-KAZE, along with various RANSAC implementations such as the original RANSAC algorithm, MAGSAC++, GC-RANSAC, DEGENSAC, etc. Both BF and FLANN matchers were evaluated, with the options to use exclusion masks and apply CLAHE. We conducted extensive grid searches to optimize parameters such as the SNN ratio threshold ($\theta_{\text{SNN}}$), the maximal number of extracted keypoints ($K$ and $K_{\text{ref}}$), the downsample ratio ($\rho$), and the epipolar threshold ($\eta$), aiming to balance accuracy and computational efficiency. Following recommendations from recent studies [127], we fixed $\tau$ at 0.999999 due to its minimal sensitivity to changes and set the maximum number of iterations $\Gamma$ at 5,000 for all RANSAC algorithms to ensure sufficient convergence. Finally, we opted for unidirectional matching, avoiding bidirectional approaches to reduce computational demands.
@@ 3.6.3 Optimization of orthophoto cut-out resolution for accurate georeferencing
Figure 11: Locations of the GCPs for the selected intersections L and P.
@@ 4.1 Traffic data extraction
The spatio-temporal diagram in Fig. 16 visualizes traffic flow dynamics at intersection L, highlighting the temporal progression of vehicle movements and queue formation and dissipation across six lanes. To maintain legibility, the diagram shows trajectories from a single video. Notable patterns include distinct queue buildup and clearance in the left-turn lane (Lane 1), consistent flow in straight-ahead lanes (Lanes 2-5), and dynamic stop-and-go behavior in the right-turn lane (Lane 6). Shorter trajectories, particularly in Lanes 1 and 2, correspond to motorcycles changing lanes or exiting the road section, causing intermittent traffic patterns. The inclusion of road section and lane-level information in our dataset enables users to create similar spatio-temporal analyses, facilitating granular studies of traffic behavior that are often limited by other datasets.
@@ Acknowledgments
This work was supported by research grants from the Swiss National Science Foundation under NCCR Automation (No. 51NF40_180545, N.G. and R.F.), Innosuisse (No. 101.645 IP-ENG, N.G. and R.F.), the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. 2022R1A2C1012380, H.Y. and H.C.), and the Open Research Data (ORD) Program of the ETH Board (R.F.). The authors express their gratitude to Sohyeong Kim, Yura Tak, and Weijiang Xiong for their invaluable assistance in the data-wrangling process, Artem Vasilev for his dedicated efforts in data annotation, Jasso Espadaler Clapés for his insightful contributions to vehicle dimension estimation, and Muhammad Ahmed for his support in creating the experimental orthophoto for validation purposes and his contributions to road segmentation. We also thank DroMii company for producing the primary orthophoto and Stanford Center at the Incheon Global Campus (SCIGC) for providing the on-board recorded trajectories of their autonomous vehicle.

### v3

@@ Abstract
Traffic monitoring , Machine vision , Vehicle tracking , Video image processing , Georeferenced vehicle trajectories , Multi-drone data collection , Urban traffic
@@ 1 Introduction
Traditional traffic monitoring methods, such as loop detectors and manual counting, are becoming obsolete in the era of smart cities due to their inflexibility and limited scope [1]. Stationary camera surveillance, although effective, presents economic challenges due to high installation costs and limited field of view (FoV), leading to blind spots. Data from global navigation satellite system (GNSS) and connected vehicles also lack the versatility needed for comprehensive traffic monitoring. Several vehicle detection and tracking technologies have emerged, including camera-based [2, 3, 4], LiDAR-based [5, 6], magnetics-based [7, 8], and radar-based [9, 10] systems. Although each offers certain advantages, they still face limitations in adaptability and coverage. Accurate estimation of traffic states allows adjustments in traffic control strategies [11], improving network utilization [12, 13] and road user experiences, as well as delivering economic and ecological benefits.
@@ 2.5 Availability of datasets
Vehicle annotation datasets, such as CARPK [110], CyCAR [111], HARPY [112], RAI4VD [113], UAVDT [114], UIT-ADrone [115], and VisDrone [116], serve as valuable training resources for object detection models. Nonetheless, they frequently utilize basic annotation methods, cover a restricted or incoherent range of vehicle classes, and exhibit varying degrees of annotation accuracy, thereby potentially affecting the reliability and consistency of DL models trained on these datasets. To mitigate these limitations, we incorporated these datasets in conjunction with our custom Songdo Vision dataset [29] and employed a multi-phase object detector training pipeline, which is described in Section 3.3, where Table 1 presents the relevant metadata for these datasets.
@@ 3.3 Object detection
The datasets used for training, comprising images and annotations from eight public datasets and our Songdo Vision dataset (detailed in Section 3.2), are summarized in Table 1, which also specifies their usage across the BASE and FINE training stages. The public datasets include a variety of aerial images, characterized by varying altitudes, camera angles, perspectives, resolutions, illumination conditions, image quality, and train/test split ratios. For the public datasets, we adhered to the original train/test splits whenever available. In cases where no such split was provided, we applied an 80/20 division to maintain uniformity. Furthermore, datasets containing a high proportion of repetitive images were proportionally down-sampled to maintain a balanced representation across sources. This ensured that no single dataset dominated the training process, enhancing the model’s ability to generalize across diverse traffic conditions.
Table 1: Vision datasets used in multi-stage object detector training.
@@
To ensure consistency with our 140m–150m BEV scenes in Songdo, we performed extensive preprocessing and filtering of these datasets. Although a comprehensive description of these steps is omitted here for brevity, we focused on consolidating the diverse class labels found in the public datasets to streamline our trajectory extraction process, which utilizes four primary vehicle categories and two auxiliary categories: $\mathcal{C}={0,,1,,2,,3,,4,,5}$, where 0 corresponds to cars (including vans and light-duty vehicles), 1 to buses, 2 to trucks, 3 to motorcycles, 4 to pedestrians, and 5 to bicycles. While only the four primary vehicle categories were used during inference, the auxiliary categories (pedestrians and bicycles) were retained during training to improve overall classification performance. To ensure that our FINE dataset meets stringent training specifications, we excluded certain datasets, such as VisDrone and UAVDT. For example, VisDrone’s annotations primarily originate from non-BEV environments at lower altitudes, resulting in inconsistent object size representations and varied camera settings that complicated integration. Additionally, this dataset exhibited mediocre annotation accuracy, further justifying its exclusion from our FINE dataset.
@@
Table 2: Annotation counts and class distributions across training and test sets for the BASE and FINE dataset groupings, with our Songdo Vision dataset also shown separately.
@@
Songdo Vision, comprising 5,419 annotated images, was split into training and test sets, with 20% (1,084 images) designated as the test set, allocated for both model validation during training and final testing of our top-performing object detector. Table 2 shows the class distribution across the sets. Our top-performing model was selected based on its highest mean Average Precision (mAP) at an IoU of 0.5 (mAP@50) across all classes. As shown in Table 3, this model demonstrated remarkable performance across all vehicle types, achieving an overall mAP@50 of 0.951 and an exceptional mAP@50 of 0.888 for motorcycles, despite their smaller pixel size. Beyond conventional detection metrics, we also assessed the positional accuracy of BB centers, revealing a mean Euclidean error of approximately 2.21 ± 1.99 pixels (12.2 ± 10.9 cm), further reinforcing the reliability of our detection step for trajectory extraction. These results highlight the model’s robustness in detecting small objects in BEV drone videos. To briefly define the key metrics: Precision indicates the proportion of correctly identified vehicles out of the total detected, while Recall reflects the ratio of correct detections to the actual number of vehicles in a class. The mAP@50 measures detection accuracy using an IoU threshold of 0.5, and mAP@50-95 evaluates performance over a range of IoU thresholds (0.5 to 0.95), with a step size of 0.05, capturing variability in BB overlaps.
@@
The confusion matrix in Fig. 4 (left) illustrates the model’s classification performance, highlighting accuracy and misclassification patterns across vehicle classes. The precision-recall curve in Fig. 4 (right) further demonstrates strong detection performance, with low false positive and false negative rates. While smaller objects like pedestrians and bicycles are occasionally misclassified as motorcycles, particularly near pedestrian crossings or bike lanes, these instances were not filtered due to the inherent challenges in accurate differentiation. Despite this, the model’s robust performance on motorcycles, though slightly lower than for larger vehicles, highlights the challenges of detecting smaller objects in drone footage.
@@ 3.5.2 Method and parameter search campaign
We implemented several prominent local feature detectors including SIFT, RSIFT, BRISK, ORB, KAZE, and A-KAZE, along with various RANSAC implementations such as the original RANSAC algorithm, MAGSAC++, GC-RANSAC, DEGENSAC, etc. Both BF and FLANN matchers were evaluated, with the options to use exclusion masks and apply CLAHE. We conducted extensive grid searches to optimize parameters such as the SNN ratio threshold ($\theta_{\text{SNN}}$), the maximal number of extracted keypoints ($K$ and $K_{\text{ref}}$), the down-sample ratio ($\rho$), and the epipolar threshold ($\eta$), aiming to balance accuracy and computational efficiency. Following recommendations from recent studies [127], we fixed $\tau$ at 0.999999 due to its minimal sensitivity to changes and set the maximum number of iterations $\Gamma$ at 5,000 for all RANSAC algorithms to ensure sufficient convergence. Finally, we opted for unidirectional matching, avoiding bidirectional approaches to reduce computational demands.
@@ 3.6.3 Optimization of orthophoto cut-out resolution for accurate georeferencing
Figure 11: Locations of the GCPs for the selected intersections L and P. Each image has been slightly dimmed to enhance the visibility of the overlaid markers.
@@ 4.1 Traffic data extraction
The spatio-temporal diagram in Fig. 16 visualizes traffic flow dynamics at intersection L, highlighting the temporal progression of vehicle movements and queue formation and dissipation across six lanes. To maintain legibility, the diagram shows trajectories from a single video. Notable patterns include distinct queue buildup and clearance in the left-turn lane (Lane 1), consistent flow in straight-ahead lanes (Lanes 2-5), and dynamic stop-and-go behavior in the right-turn lane (Lane 6). Shorter trajectories, particularly in Lanes 1 and 2, correspond to motorcycles changing lanes or exiting the road section, causing intermittent traffic patterns. The inclusion of road section and lane-level information in our dataset enables users to create similar spatio-temporal analyses, facilitating granular studies of traffic behavior that are often limited by other datasets. A video demonstration of our extraction pipeline [30], Geo-trax, is available at https://youtu.be/gOGivL9FFLk.
@@ 5 Conclusion
CRediT authorship contribution statement
Robert Fonod: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Visualization, Writing – original draft, Writing – review & editing, Visualization, Supervision, Project administration. Haechan Cho: Methodology, Software, Validation, Investigation, Data curation, Visualization. Hwasoo Yeo: Conceptualization, Resources, Funding acquisition, Supervision, Project administration, Funding acquisition. Nikolas Geroliminis: Conceptualization, Investigation, Resources, Writing – review & editing, Supervision, Project administration, Funding acquisition.
@@ Acknowledgments
This work was supported by research grants from the Swiss National Science Foundation under NCCR Automation (No. 51NF40_180545, N.G. and R.F.), Innosuisse (No. 101.645 IP-ENG, N.G. and R.F.), the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. 2022R1A2C1012380, H.Y. and H.C.), and the Open Research Data (ORD) Program of the ETH Board (R.F.). The authors express their gratitude to Sohyeong Kim, Yura Tak, and Weijiang Xiong from EPFL for their invaluable assistance in the data-wrangling process, Artem Vasilev for his dedicated efforts in data annotation, Jasso Espadaler Clapés for his insightful contributions to vehicle dimension estimation, and Muhammad Ahmed for his support in creating the experimental orthophoto for validation purposes and his contributions to road segmentation. We also thank DroMii company for producing the primary orthophoto and Stanford Center at the Incheon Global Campus (SCIGC) for providing the on-board recorded trajectories of their autonomous vehicle. We also thank the research teams of Prof. Simon Oh (Korea University) and Prof. Minju Park (Hannam University) for their assistance during the data collection campaign, including the provision of drone equipment and student support.